# Seed-Music: A Unified Framework for High Quality and Controlled Music Generation

### v1

@@ 3.3 Vocoder Latent-based Pipeline
Figure 4: Seed-Music pipeline with vocoder latents as intermediate representation. (1) Various input types are fed into DiT via cross-attention, prefix, or time-aligned conditioning. (2) The diffusion transformer model predicts the continuous vocoder latents. (3) The acoustic vocoder then produces high-quality 44.1kHz stereo audio.
@@
  • Time-aligned inputs that span multiple tracks: Temporal signals such as melody contours, intensity curves, and instrumental stems that are time-aligned conditioning inputs can be added at each step of the denoising process.
  • Multi-channel outputs: Supported when multi-channel output examples are provided during training. For instance, the model can generate multiple musically distinct stems (e.g., vocals, bass, drums, and guitar), enabling downstream production scenarios like mashups and remixing. These stem-level training examples can be sourced from Music Source Separation (MSS) models.
@@ 3.4 Model Training and Inference
Inspired by these findings, we explore RL methods (Ahmadian et al., 2024; Prabhavalkar et al., 2018; Wang et al., 2024b; Sutton et al., 1999; Schulman et al., 2017) to improve the alignment of generated output with various input control signals and enhance musicality. Reward models we considered include: the edit distance between the original lyrics prompt and the lyrics transcription extracted from the generated audio, the genre prediction accuracy by comparing the input genre with the detected genre of the audio output, and the match between a song structure prompt and the detected structure in the generated audio. Additional reward models based on tempo, instrumentation, audio references, and singing voice references can be used to dictate what musical attributes are emphasized in the generation output. Moreover, incorporating human feedback (Ouyang et al., 2022) can produce reward models that capture subtle user preferences beyond the above objective metrics. We leave the thorough study of RL for future work.
@@ 4.1 Lyrics2Song
Lyrics2Song generates vocal music performances conditioned on user-provided lyrics and music style descriptions. This task utilizes the audio token-based pipeline (see Section 3.1), leveraging tokenization and auto-regressive techniques to align multi-modal data (i.e., lyrics, style tags, and audio) and enable streaming decoding for fast, responsive interactions.
@@
This system supports both cohesive short-form audio clip generation [^1] https://team.doubao.com/seed-music/shortform-audio-generation and full-length track production [^2] https://team.doubao.com/seed-music/longform-audio-generation. The generated audio showcases expressive and dynamic vocal performances with engaging melodies as well as instrumentals that span a wide variety of genres and moods, delivering a mature level of musicality.
Vocal Music Generation with Audio Reference
@@
In addition to style descriptions, our system also supports audio inputs to guide the music styles. The listening examples [^3] https://team.doubao.com/seed-music/audio-prompting demonstrate how the outputs are generated by referencing the musical styles in the audio prompts. As describing desired music with text or tags is usually less straightforward for novice users, we believe using the audio prompts can better convey the intrinsic needs. Our system supports two modes of audio prompting. In “continuation mode”, the tokenizer extracts the sequence of audio tokens from the audio is used as the prefix, preserving more details of the reference and providing a stronger supervising signal to generation. In “remix mode”, the input audio is extracted into an embedding vector in a pre-trained joint embedding space of text and audio (Huang et al., 2022), so that it can be concatenated in the prefix for the auto-regressive LM. The embedding vector summarizes the global characteristics of the audio prompt.
In both modes, we demonstrate the versatility of audio prompting with pairing different lyrics inputs for generation and the lyrics from the reference song. When the lyrics used for generation closely follow the original lyrics in the reference song, we observe seamless continuation of the reference, with echoing melody line and song structure. When the new lyrics differ significantly from the reference song in style—such as syllable count, structure, or language—the similarity between the generated result and the reference audio weakens. Despite this, there remains a natural resemblance in musical motifs, rhythm patterns, instrumentation, voice timbre, and overall vibe.
Instrumental Music Generation The same pipeline supports pure instrumental music generation, where the creator omits the lyrics and simply provides descriptors for the desired music. We provide audio samples [^4] https://team.doubao.com/seed-music/instrumental-music-generation in a wide range of styles with creative musical transitions to fit precise timing prompts for section transitions.
@@
Evaluation Metrics During model development, we considered the following quantitative metrics, which are also repurposed as reward models in reinforcement learning of the audio token based LM.
  • WER: We employ the Whisper-large-v3 model for English vocal music and the Paraformer-zh for Chinese vocal music to transcribe the generation example and compute the word (Pingyin) error rate and against the lyrics prompt. Although WER is a typical metric for TTS, it is not a perfect measure for vocal music quality. Sung words usually introduce elongated vowels and consonants as well as larger ranges in pitch and non speech-like cadences. These are attributes that an automatic WER could penalize but are critical for dynamic and expressive vocal music.
  • MIR classification: To measure the adherence between generated audio and style prompt, we use in-house Music Information Retrieval (MIR) models to run inference on the generated audio, and check the correspondence between the predicted genre, mood, vocal timbre, song structure with timing information, against the music elements mentioned in the style prompt.
@@
For qualitative metrics, we use the Comparative Mean Opinion Score (CMOS) based on a team of musically-trained raters. We define the following critical dimensions for assessment.
@@
  • Musicality measures musically semantic attributes and was decomposed into interesting vocal melodies, appropriate use of harmony, presence of idiomatic musical forms like theme and variation, plausible structure, suitable chord progressions, characteristic rhythmic patterns and full-bodied instrumentation arrangement.
@@
  • Audio Quality measures acoustic attributes including the clarity of vocals, realism of instruments, detail across the frequency spectrum and transients of drums and onsets. Raters also consider unwanted audio artifacts like distortion, muffled audio or missing frequency bands.
@@
  • Prompt Adherence measures how closely the generated audio matches the lyrics input and other style prompts.
Unlike in speech domains, where Word Error Rate (WER) and Automatic Speaker Verification (ASV) are widely used to compare TTS and ASR models, no such straightforward quantitative metric or benchmark datasets exist in music. Moreover, musicality, the most important aspect, is nearly impossible to be quantified by an objective metric. We hope our documented methodology can guide subsequent research in vocal music generation. We note the challenge in presenting quantitative and qualitative statistics that can be compared against for music generation, and thus we encourage readers to directly listen to the many audio demos on the website to judge the quality of our system.
@@
We have also conducted Lyrics2Song experiments with a diffusion-based pipeline as outlined in Section 3.3, and achieved on-par end-to-end performance. Nevertheless, we find the LM-based pipeline to be naturally suited for interactive use cases. Its causal nature enables streaming solutions like those described in Section 3.1 to provide a near real-time experience and allows future integration with multi-modal models (Liu et al., 2023).
@@ 4.2 Lyrics2Leadsheet2Song
In this set of experiments, we develop solutions around the system pipeline depicted in Section 3.2 to expose lead sheet tokens for musicians to interact with as an intermediate step in the end-to-end music audio generation process.
Lyrics2Leadsheet We developed a rule-based symbolic music codec based on (Chen et al., 2024c) to encode the symbolic features of a music audio piece into a sequence of lead sheet tokens. In Figure 5, we visualize how the information of different tracks, including lyrics, vocal melody, piano, guitar, drum, chord, are encoded in the token sequence. We interleave different tracks on a bar-by-bar basis; and for individual element (e.g. phoneme, note and chord) in each track, we encode the onset, duration, and pitch information if available.
Figure 5: Illustration of the REMI-style symbolic music encoding.
We show with the demo examples [^5] https://team.doubao.com/seed-music/lyrics-to-leadsheet how the LM Generator, trained with the transcribed lead sheets with the MIR models discussed in Section 3.2, learns to predict phoneme-aligned notes and genre-appropriate melodies and rhythms based on the lyrics and style prompts.
Leadsheet2Song Leadsheet2Song generalizes the task of transforming symbolic representations into full mixes of music audio. By inspecting the listening examples [^6] https://team.doubao.com/seed-music/leadsheet-to-song, one can hear how the generated audio follows vocal melodies, phonemes, instrumental melodies, rhythms and chords specified in the lead sheet tokens, and renders the full mixture of natural and expressive music performance.
Leadsheet2Vocals (Singing Voice Synthesis) Instead of outputting a full mix, the lead sheet token based generative system can be configured to generate individual stems. Singing Voice Synthesis (SVS) becomes a straightforward subset of lead sheet token-based pipelines as shown in the demo [^7] https://team.doubao.com/seed-music/leadsheet-to-vocals.
@@ 4.3 Music Editing
In this section, we consider the task of making edits to music audio as a post-production process. The non-causal nature of diffusion-based approach proposed in Section 3.3 makes it naturally suited for editing tasks. For example, in text-conditioned in-painting, a diffusion module’s visibility of context before and after the masked audio segments guarantees smoother transitions (Wang et al., 2023c). We formulate the problem as a lead sheet conditioned in-painting task to train the Diffusion Tranformer model. Upon inference, the modified lead sheet is used as conditioning input, and the segment of audio corresponding to the modified part of the lead sheet is masked and re-generated.
The listening examples [^8] https://team.doubao.com/seed-music/editing-lyrics show the system’s ability to precisely alter the sung lyrics while keeping the melody and backing track unchanged. Conversely, creators can also precisely alter the melody while keeping the melody and backing track unchanged [^9] https://team.doubao.com/seed-music/editing-melody. We demonstrate this for English and Mandarin Chinese. We are particularly excited by this new form of “generative audio editing”, which retains the natural performance and musical semantics of the original track. These are post-production edits that historically would have required original project stems and MIDI tracks.
4.4 Singing Voice Conversion
In the context of Seed-Music, singing voice conversion (SVC) is the final piece of the suite that bridges a creator’s own voice with the generative systems. Although our VC system is similar to Seed-TTS, vocal music VC presents new challenges distinct from speech VC (Arik et al., 2018; Anastassiou et al., 2024).
  • Vocal mixture: Vocal music data natively consists of vocal track and background music, whereas speech data has no background music. Although modern MSS algorithms can be applied and the resulting vocal stem treated with conventional voice conversion techniques, this approach is not immune from the artifacts introduced in the MSS process. Our goal was to develop a scalable system that could directly operate on the mix of vocal track and background music.
@@ 4.4 Singing Voice Conversion
  • Vocal range: The range of vocal singing is much wider than vocal speech. In the context of zero-shot singing voice conversion, the system must robustly extrapolate the synthesized singing voice to fundamental frequencies beyond the reference vocal.
@@
  • Vocal technique: Vocal singing techniques are much more complex than speech. The same vocalist can sound very different when singing in operatic mezzo-soprano style, musical theater belting style or jazz scatting style. Singing VC has to deliver an expressive rendition of these techniques on top of regular speech attributes like clear pronunciation and prosody.
  • Vocal vs speech reference: In speech VC, it is a straightforward task to obtain reference speech from an end-user. In singing VC however, it is non-trivial for the end-user to provide a reference singing vocal performance. Since our goal was to creatively empower normal individuals, our system was deliberately designed to be no-singing-required. It can produce vocal music outputs given only a short speech clip as reference.
  • Amateur vs professional singing: Speech benefits from a relatively easy procedure for finding paired data between an input speaker and target speaker e.g. recording different voices reading the same passage. However, no such large scale data exists for modeling an amateur singer and professional singer. A vocal music VC system must translate non-professional singing into an improved performance, e.g. if the end-user provides out-of-tune singing, the VC result should retain the tone of voice but actually be in tune.
Naturally, when the input reference vocal and target conversion vocal are similar, i.e. both are male voices singing in English, the results are the best. Other combinations are more challenging, such as when the input reference vocal is a female speaking (not singing) in Mandarin and the target conversion vocal is a female singing in Mandarin or even harder still, singing in English. The listening examples [^10] https://team.doubao.com/seed-music/singing-voice-conversion demonstrate how our VC system is able to handle these combinations.
@@ 5 Conclusion
In this report, we introduced Seed-Music, a family of generative models for music creation tasks. We demonstrated the generation of high quality vocal music conditioned on natural language style description, or audio reference, or music score. We proposed an interpretable lead sheet tokenizer, which empowered precise audio editing workflows. Our approach revolves around three intermediate representations, each chosen to best support musical workflows from ideation to generation and editing. The flexible nature of the framework opens the door to many exciting future works, including more granular stem generation and multi-modal conditioning signals.
Seed-Music lowers the barriers to artistic creation and musical expression. We strongly believe the demos shown in this report can empower creators across the spectrum of novices and professionals. For example, our coupling of text-to-music pipelines with zero-shot singing voice conversion was intended to integrate novices more deeply into the creative process. Rather than manipulating music from a distance, novices bring their own unique voices and identities to participate in creative ideation.
Music is also integral to many complementary mediums including shortform video, longform cinema, games and AR/VR experiences. Real-time conditioning and rendering of generative music opens up entirely new interactions beyond traditional playback of audio files. We imagine entirely new artistic mediums where generative music responds to conditioning signals beyond text but in-game storylines and visual art-style.
For professionals, our lead sheet tokens were engineered from the ground-up to respect and integrate into workflows of musicians, composers, vocalists and artists. We believe lead sheet tokens can evolve to become the symbolic standard for music language models, similar to today’s industry-standard MIDI for traditional music making. Musicians and producers would be able to leverage the power of generative models through familiar means of control over melodic and rhythmic elements. Moreover, the ability to edit and manipulate recorded music in ways that preserve music semantics is a meaningful time-saving and cost-saving capability across the industry. We are also excited by future stem-based generation and editing workflows beyond the vocal track. Professionals can audition musical ideas with greater speed and experience higher chances of landing on “happy accidents” crucial to the creative process.

### v2

@@ 3.3 Vocoder Latent-based Pipeline
Figure 4: Seed-Music pipeline with vocoder latents as intermediate representation. (1) Various input types are fed into DiT via cross-attention, prefix, or temporal conditioning. (2) The diffusion transformer model predicts the continuous vocoder latents. (3) The acoustic vocoder then produces high-quality 44.1kHz stereo audio.
@@
  • Temporal inputs that span multiple tracks: Time-varying signals such as melody contours, intensity curves, and instrumental stems that are time-aligned conditioning inputs can be added at each step of the denoising process.
  • Multi-channel outputs: Supported when multi-channel output examples are provided during training. For instance, the model can generate multiple musically distinct stems (e.g., vocals, bass, drums, and guitar), enabling downstream production scenarios like mashups and remixing. These stem-level training examples can be obtained from Music Source Separation (MSS).
@@ 3.4 Model Training and Inference
Inspired by these findings, we explore RL methods (Ahmadian et al., 2024; Prabhavalkar et al., 2018; Wang et al., 2024b; Sutton et al., 1999; Schulman et al., 2017) to improve the alignment of generated output with various input control signals and enhance musicality. Reward models we considered include: the edit distance between the original lyrics prompt and the lyrics transcription extracted from the generated audio, the genre prediction accuracy by comparing the input genre with the detected genre of the audio output, and the match between a song structure prompt and the detected structure in the generated audio. Additional reward models based on tempo, instrumentation, audio references, and user voice prompts can be used to dictate what musical attributes are emphasized in the generation output. Moreover, incorporating human feedback (Ouyang et al., 2022) can produce reward models that capture subtle user preferences beyond the above objective metrics. We leave the thorough study of RL for future work.
@@ 4.1 Lyrics2Song
Lyrics2Song generates vocal music performance conditioned on user-provided music style descriptions and lyrics with section tags (e.g., ‘verse’, ‘chorus’, and ‘bridge’) (wang2022musfa). This task utilizes the audio token-based pipeline (see Section 3.1), leveraging tokenization and auto-regressive techniques to align multi-modal data (i.e., lyrics, styles, tags, and audio) and enable streaming decoding for fast, responsive interactions.
@@
This system supports both cohesive short-form audio clip generation [^1] https://team.doubao.com/seed-music/shortform-audio-generation and full-length track production [^2] https://team.doubao.com/seed-music/longform-audio-generation. The generated audio showcases expressive and dynamic vocal performances with engaging melodies as well as instrumentals that span a wide variety of instrumentation and genres, delivering a mature level of musicality.
Vocal music generation with audio reference.
@@
In addition to style descriptions, our system also supports audio input as a prompt to guide music generation. The listening examples [^3] https://team.doubao.com/seed-music/audio-prompting demonstrate how outputs are generated by referencing the musical styles of the audio prompts. Since describing desired music with text or tags can be less intuitive for novice users, audio prompts provide a more effective way to communicate musical intent.
Our system supports two modes of audio prompting: continuation mode and remix mode. In continuation mode, audio tokens extracted from the audio reference are concatenated in the prefix to continue auto-regressive generation, ensuring strong structural, melodic, and sonic similarities to the reference. In remix mode, the audio reference is converted into an embedding vector within a pre-trained joint text-audio embedding space (Huang et al., 2022). This embedding, which summarizes the global characteristics of the audio reference, is then incorporated into the prefix to guide the generation, allowing the generated audio to adopt different styles.
In both modes, our model demonstrates a strong ability to maintain coherence between the input lyrics and the inherent lyrics in the audio reference, even without the aid of automatic lyrics transcription. When the input lyrics are structurally and semantically similar to those in the audio reference, the model tends to resample the melody and structure from the reference. However, when the input lyrics differ significantly in style (e.g., language, structure, rhyme), the coherence weakens. Despite this, the model effectively maintains natural rhythm patterns, instrumentation, vocal quality, and overall musical motifs.
Instrumental music generation. Although the audio token-based pipeline is primarily designed for vocal music generation, it also supports instrumental music generation if the lyrics input contains only section tags without text. We provide several generated instrumental examples [^4] https://team.doubao.com/seed-music/instrumental-music-generation in a wide variety of styles, with each section unfolding at the specified time and demonstrating clear structural transitions between sections (e.g., verse to chorus).
@@
Evaluation metrics. We used the following quantitative metrics to assess generation quality during development. These metrics were also repurposed as reward models in the Reinforcement Learning process for the auto-regressive LM.
  • Word error rate (WER): We use in-house singing lyrics transcription model, which supports both English and Mandarin Chinese languages, to transcribe the generated audio and compute the word (or Pinyin) error rate relative to the lyrics prompt. While useful, WER is not a perfect measure of vocal quality in music due to factors like elongated vowels, consonants, pitch variations, and non-speech-like rhythms in sung words. These characteristics can introduce noise into the WER calculation.
  • Music tagging performance: To evaluate the alignment between the generated audio and input style descriptions, we use in-house music tagging and structural segmentation models to predict high-level musical attributes from the generated audio, including genre, mood, vocal timbre, vocal gender, and structural sections. These predicted attributes are then compared to the input style descriptions and section tags, with average precision scores serving as the quantitative metric for relevance.
@@
For qualitative evaluation, we use the Comparative Mean Opinion Score (CMOS), based on feedback from a team of musically-trained raters. We define the following three dimensions for assessment:
@@
  • Musicality evaluates musical attributes, including novelty of vocal melodies, appropriate use of harmony, idiomatic musical forms (e.g. theme, variation), coherent structure, suitable chord progressions, characteristic rhythmic patterns, and well-rounded instrumentation.
@@
  • Audio quality assesses acoustic characteristics such as vocal clarity, instrument realism, detail across the frequency spectrum, and the sharpness of drum transients and onsets. Raters also consider any unwanted audio artifacts, such as distortion, muffling, or missing energy in certain frequency bands.
@@
  • Prompt adherence measures how closely the generated audio aligns with the input lyrics and style prompts.
In the speech domain, benchmark datasets are established to evaluate TTS systems using metrics like WER and Automatic Speaker Verification (ASV). However, there are no equivalent benchmarks for music generation that provide quantitative scores. Additionally, it is important to note that musicality—a key factor in assessing music generation quality—is highly subjective and challenging to quantify with objective metrics. We encourage readers to listen to the provided audio demos to better assess the quality of our system.
@@
Audio tokens versus vocoder latents. We also conducted Lyrics2Song experiments using the vocoder latent-based pipeline (see Section 3.3), achieving performance comparable to the audio token-based pipeline (see Section 3.1). However, we find the auto-regressive LM to be inherently better suited for interactive applications than the diffusion model. Its causal architecture enables streaming solutions that provide a near real-time experience, while also allowing for future integration with multi-modal models (Liu et al., 2023).
@@ 4.2 Lyrics2Leadsheet2Song
The Lyrics2Leadsheet2Song pipeline is a two-step process for achieving the Lyrics2Song task: Lyrics2Leadsheet and Leadsheet2Song. In the first step, lead sheet tokens are generated from the input lyrics and style descriptions. In the second step, music audio is produced from the lead sheet tokens. The overall pipeline is illustrated in Section 3.2. Lead sheet tokens enable user involvement in the generation process by allowing edits to melody, chords, instrumentation, and tempo before the final rendering.
Lyrics2Leadsheet. We developed a rule-based symbolic music encoding scheme based on (Chen et al., 2024c) to encode the symbolic features of a music audio piece into a sequence of lead sheet tokens. As illustrated in Figure 5, the scheme encodes lyrics and various musical events. It recognizes eight types of events: lyric phoneme, bar, chord, vocal note, bass note, piano note, guitar note, and drum note. Each type of events, except for ‘bar’, is represented as a distinct “track” in the lead sheet tokens. Bar events define the basic temporal structure, with tracks interleaved on a bar-by-bar basis. For each event (e.g., phoneme, note, chord) within a track, we encode onset, duration, and pitch values when applicable.
Figure 5: Illustration of the REMI-style symbolic music encoding scheme.
As described in Section 3.2, the Generator is trained using lead sheet tokens extracted by our in-house MIR models. Here are some examples [^5] https://team.doubao.com/seed-music/lyrics-to-leadsheet demonstrating how the auto-regressive LM predicts phoneme-aligned notes with genre-appropriate melodies and rhythms according to the input lyrics and style prompts.
Leadsheet2Song. The pipeline of Leadsheet2Song involves rendering a full audio mix from a sequence of lead sheet tokens. In the demo examples [^6] https://team.doubao.com/seed-music/leadsheet-to-song, we showcase how the generated vocal music follows the vocal melodies, phonemes, rhythms, chord progressions, and instrumental notes from the given lead sheet tokens. The Renderer effectively generates the nuances of natural and expressive music performance across multiple instruments, providing professionals with a powerful tool to quickly review audio results without the need for meticulous parameter adjustments in synthesizers.
Leadsheet2Vocals for singing voice synthesis. Rather than producing a full audio mix, the Lyrics2Leadsheet2Song system can be configured to generate individual stems, including vocals, drums, bass, piano, and guitar, in both symbolic notation as well as rendered audio. Singing Voice Synthesis (SVS) is one application of this system, where the model is set to output only vocal stem, as demonstrated in these examples [^7] https://team.doubao.com/seed-music/leadsheet-to-vocals.
@@ 4.3 Music Editing
In this section, we explore music audio editing as a post-production process. The non-casual nature of the diffusion-based approach, described in Section 3.3, makes it particularly well-suited for such tasks. For example, in text-conditioned in-painting, a diffusion model’s ability to access context before and after masked audio segments ensures smoother transitions (Wang et al., 2023c). We frame this as a lead sheet-conditioned in-painting task to train the DiT model. During inference, the modified lead sheet serves as conditioning input, and the audio segment corresponding to the altered part of the lead sheet is masked and re-generated.
In these listening examples [^8] https://team.doubao.com/seed-music/editing-lyrics, we demonstrate the system’s ability to precisely modify sung lyrics while preserving the melody and backing tracks using both English and Mandarin Chinese singing voices. In some cases, the lyrics are altered within the same language, while in others, the system allows for alternating between languages. Moreover, in these examples [^9] https://team.doubao.com/seed-music/editing-melody, we demonstrate how users can precisely adjust the melody within a specified time segment, while keeping the lyrics, the rest of the melody, and the backing track unchanged. This new “generative audio editing” paradigm excites us, as it retains the musical performance and essential qualities of the original track—something that was previously complex or nearly impossible without re-recording the vocals along with the original instrumental stems.
4.4 Zero-shot Singing Voice Conversion
One of the most intuitive ways for creators to edit vocal music is by converting the vocal timbre to match their own voice. This section explores the singing Voice Conversion (VC) system as the final component of the Seed-Music suite. Although our singing VC method shares similarities with the speech VC introduced in Seed-TTS (Anastassiou et al., 2024), voice cloning and conversion in the context of vocal music generation present greater challenges compared to the speech domain (Arik et al., 2018):
  • Vocal mixture: Vocal music typically consists of vocals and background instrumental tracks, with strong coherence between them in terms of harmony and rhythm. In contrast, speech signals often contain background environmental sounds unrelated to the speech content. While modern MSS models can isolate vocals, they often introduce artifacts that degrade quality. Our goal is to develop a scalable system capable of directly processing the mixture of vocals and background tracks without relying on MSS, thereby avoiding these artifacts.
@@ 4.4 Zero-shot Singing Voice Conversion
  • Vocal range: The pitch range of singing voices is much wider than that of speech. For zero-shot singing VC, the system must generalize the pitch range of the reference voices to the synthesized singing voices, requiring strong robustness from the model.
@@
  • Vocal technique: Singing voices are highly expressive and involve far more techniques than speech. The same vocalist can sound vastly different when performing in styles like operatic mezzo-soprano, musical theater belting, or jazz scatting. A singing VC system must accurately capture and reproduce these expressive techniques, in addition to handling regular speech attributes like clear pronunciation and prosody.
  • Singing versus speech reference: In VC applications, users typically provide speech as their reference voice, whether for speech or singing voice synthesis. Our system is specifically designed to accept reference voices regardless of whether they are speech or singing. It can effectively perform singing VC using a short speech clip as the reference.
  • Amateur versus professional singing: There is significantly less paired data available for amateur versus professional singing compared to speech VC data. This makes singing VC particularly challenging, as the model has to adapt to non-professional singing inputs and convert them into professional-quality performances. For instance, if users provide singing references that are out of tune, the singing VC system must not only capture the tone of their voices but also correct the pitch.
The listening examples [^10] https://team.doubao.com/seed-music/singing-voice-conversion demonstrate how our singing VC system performs across different scenarios. The quality of the outcomes depends largely on the similarity between the reference voices and the target singing signals. For instance, when both are male voices singing in English, the results are optimal. However, handling cross-gender and cross-language cases is more challenging, often leading to issues such as artifacts, distortions, and inconsistencies in pronunciation.
@@ 5 Conclusion
In this report, we have introduced Seed-Music, a comprehensive suite of music generation and editing systems designed to support diverse music creation workflows. We have also demonstrated how the system generates high-quality vocal music conditioned on multi-modal inputs, including lyrics, style descriptions, audio references, music scores, and voice prompts. Our unified framework addresses various use cases using three intermediate representations (i.e., audio tokens, lead sheet tokens, and vocoder latents) and their associated pipelines, providing users with flexible tools to move from ideation to generation and editing.
From application perspectives, Seed-Music lowers the barriers to artistic creation and musical expression. We believe the demos shown in this report can empower a wide range of creators, from novices to professionals. For example, the integration of text-to-music system with zero-shot singing voice conversion allows novices to engage more deeply in the creative process. Instead of interacting with music from a distance, novices can bring their own unique voices and identities into the process, enhancing creative ideation.
Music is also a key component of complementary media such as short-form videos, films, games, and AR/VR experiences. For more applications, real-time conditioning and rendering of generative music introduce entirely new forms of interaction, going beyond traditional audio playback. We envision new artistic mediums where generative music responds to conditioning signals not only from text but also from in-game narratives and visual art styles.
For professionals, the proposed lead sheet tokens are designed to integrate seamlessly into the workflows of musicians, composers, vocalists, and artists. We believe these tokens have the potential to evolve into a symbolic standard for music language models, much like MIDI has for traditional music production. Musicians and producers could harness the power of generative models while maintaining familiar control over melodic, harmonic, and rhythmic elements. Furthermore, the ability to edit and manipulate recorded music while preserving its musical semantics offers significant time and cost savings for the industry. We are particularly excited about future developments in stem-based generation and editing, which would extend beyond vocal tracks. These capabilities will allow professionals to explore musical ideas more efficiently, increasing the likelihood of discovering the “happy accidents” that are often pivotal to the creative process.